A random sample of 600 high school girls showed that during the previous month 42 took diet pills to lose weight. The researcher wants to test the hypothesis that less than 10% of high school girls take diet pills to lose weight. The test statistic is
Answer by stanbon(57424)

A random sample of 600 high school girls showed that during the previous month 42 took diet pills to lose weight. The researcher wants to test the hypothesis that less than 10% of high school girls take diet pills to lose weight.
The test statistic is:
z(42/600) = (0.07-0.10)/sqrt{0.1*0.9/600] = -2.4495
=====================
Cheers,
Stan H.
=============